# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### v2

@@ 6 Evaluation
Figure 5: We evaluate GPT-3.5 Turbo (GPT-3.5), GPT-4 Turbo revision 1106 (GPT-4), Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs). We denote the similarity score such that higher is better. The similarity score is measured according to a reference baseline and normalized to exclude noise.
@@ 7 Discussion
In this section, we address the limitations of SymbolicAI and the future directions we are focusing on. Some of the limitations stem from the inherent constraints of current technologies and dependencies on third-party systems. Additionally, the nuanced complexities of working with generative models presents further challenges.
7.1 Limitation
@@
Our empirical measure is limited by the expressiveness of the embedding model and how well it captures the nuances in similarities between two representations. Furthermore, the obtained similarity scores are highly non-linear and difficult to interpret. For instance, two representations may address the same topic, such as the problem description and its respective solution, however, when measuring their similarity we obtain similarity scores of $\sim 70\%$. We normalize this by subtracting an inherent baseline and randomness effect, however, to ensure a more holistic and robust measurement we would need a significantly larger amount of baselines and experiments. Since we were very limited in the availability of development resources, and some presented models are only addressable through costly API walls. We are actively seeking investors and supporters to scale our solution and offer a more compelling benchmark suite in the future.
@@
This research is also limited by current LLM’s capacity for reasoning and generalization. Although progress has been made, models are still prone to hallucinations and reasoning errors, especially when dealing with abstract, novel, or highly complex problem statements (Marcus, 2020). Furthermore, our framework’s reliance on the model’s ability to grasp the semantics of operations can be influenced by the training data and the model’s innate biases and conceptual understanding (Mahowald et al., 2023). We also point out that the initial development of SymbolicAI started with the GPT family of models, and we may encounter innate biases in prompt design and expressiveness when using other reference models. However, we also point out that prompt engineering instruction-based statements is not a reliable direction for improvement. We instead advocate for enhancing the resilience of models through fault tolerance, focusing on their ability to better follow semantic instructions, not syntactic idiosyncrasies. Another concern is how to assess the disentanglement of evaluations of models on downstream tasks, to avoid testing on training samples, especially for closed-source solutions like GPT.
@@ 7.2 Future Work
The goal for Algorithm 1 is to be utilized by an advanced learning agent. This agent, employing reinforcement learning methodologies (Ouyang et al., 2022; Li et al., 2023; Rafailov et al., 2023), could leverage our evaluation measure in existing benchmarks (Milani et al., 2020; Swazinna et al., 2022; Schweighofer et al., 2022) as a means to obtain reward signals to addresses a central problem in reinforcement learning, namely credit assignment (Sutton, 1984; Arjona-Medina et al., 2019; Holzleitner et al., 2020; Patil et al., 2020; Widrich et al., 2021; Dinu et al., 2022). Over time, it aims to develop the ability to autonomously generate its own plans, efficiently schedule tasks and subtasks, and carefully select the most suitable tools for each task. Our protocol lays the groundwork for this agent to learn and expand its base set of capabilities (Amaro et al., 2023), moving towards more sophisticated, self-referential orchestration of multi-step tasks. We’ve already noticed that research is shifting towards this type of methodology (Yuan et al., 2024). Furthermore, in Section 6 we’ve only considered a sequential scheduler. However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution. In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a, b; Lahlou et al., 2023).
7.3 Broader Impact
With LLMs becoming more and more accessible, progress recently made possible by the vast open source contributions from Köpf et al. (2023); Touvron et al. (2023); Taori et al. (2023); Xu et al. (2023); Geng et al. (2023); Biderman et al. (2023), embedded accelerators for LLMs – or more generally NeSY engines – will be ubiquitous in future computation platforms, such as wearables, smartphones, tablets, consoles, or notebooks. Although current execution cycles are slow and error-prone, we expect to see further performance gains through improved operating system level optimizations, dedicated GPU-centric hardware refinement, and improved software interoperability. We believe that modern programming paradigms should natively support probabilistic concepts and provide a boilerplate-free set of features for constructing and evaluating generative computational graphs. This includes but is not limited to compositional, parallelizable, and simulation-based executions with polymorphic and self-referential structures. Current programming languages often have disjointed or makeshift solutions for these concepts in the context of generative processes. We believe integral probabilistic support for these concepts into modern software and hardware will unlock new programming paradigms that can fully take advantage of generative architectures. We hope the community will consider these ideas as essential components of contemporary computing.
Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure. Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs. This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time. We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023). We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing General AI agents.
7.4 Conclusion
@@
In this work, we introduced SymbolicAI, a framework that unifies generative models with an array of solvers, blending the strengths of symbolic and sub-symbolic AI paradigms within a cohesive NeSy framework. SymbolicAI equips researchers and practitioners with a comprehensive toolkit to develop contextualized and explainable NeSy AI systems capable of addressing diverse challenges effectively. We also introduce a quality measure and a benchmark for comparing and evaluating a wide range of computational tasks. SymbolicAI provides a basis for further research in advanced program synthesis, hierarchical computational graphs, the development of self-referential systems, and the integration of probabilistic methods with AI design for creating autonomous agents. Furthermore, our commitment to open-source ideals attempts to promote a culture of collaborative growth and innovation.
@@ Acknowledgement
Appendix
Appendix A Connection between Fréchet Distance and Maximum Mean Discrepancy
@@
In the next section we want to further elaobarate on the MMD and a possible score, that can be derived from it.
A.1 Extended Simplification of the MMD Calculation
@@
Appendix B Broad AI and Neuro-Symbolic Systems
Our work focuses on broad artificial intelligence (AI) (Hochreiter, 2022) (see Figure 6) through the integration of symbolic and sub-symbolic AI methodologies. Broad AI extends beyond restricted focus on single-task performance of narrow AI. In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023), as illustrated in Figure 7. This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.
Appendix C Framework Structure
@@
Type inheritance in SymbolicAI is leveraged to create new expressions, which are specialized forms of Symbol objects designed to represent parts of a computational graph. Expressions extend the capabilities of Symbol objects by providing a structured way to define complex functionalities that can later be evaluated to produce new Symbol objects or modify existing ones. In SymbolicAI, an Expression object inherits the properties of Symbol objects while also being able to define its own unique behavior through the use of a forward method, which is analogous to a computational graph node’s evaluation function. Figure 9 gives an overview of an exemplary inheritance branch. Each Expression must feature a forward method, which must be overwritten to define its behavior. The inherited __call__ method invokes the forward method, evaluating the expression and returning the result. This design pattern facilitates lazy evaluation of expressions, allowing for complex composition of expressions.
@@
The use of decorators brings several advantages (Beazley, 2009; Martelli et al., 2005; Summerfield, 2010; Lutz, 2013):
@@
  • Adaptability: Using decorators we can easily modify or extend the behavior of operations without changing their core implementation. This flexibility facilitates the generation of adaptive workflows and reliable fallback mechanisms when experimental implementations do not fulfill required constraints.
@@
In practice, consider the Expression class, which extends the functionality of ‘Symbol‘. When composing a Sequence of Expression objects, we are effectively composing operations in a predetermined order.
For instance, an expression such as:
@@ 5  Compose(),
Generating entire workflows
SymbolicAI allows the creation of entire workflows, such as writing scientific papers. The following example defines a Paper expression that takes in a sequence of expressions which are executed in sequentially. The Method expression contains a Source expression, which addresses the code base of the actual method of the paper. The RelatedWork expression contains a sequence of Cite expressions which are executed in parallel and are used to define the context of the related work section. Each expression in the sequence of expressions from Paper takes in the context of its predecessors. All expressions also link their results to a global linker object, which is used after the execution to retrieve individual results from distinct expression nodes. In Figure 10 we show the resulting computational graph of the Paper expression.
1# define the computational graph
2expression = Paper(
3  Method(
4    # link to original code base where the main method is defined
5    Source(file_link=’/path/to/.../file’),
6  ),
7  # gather resources and write the related work
8  RelatedWork(
9    Cite(file_link=’/path/to/.../file’),
10    Cite(file_link=’/path/to/.../file’),
11    ...
12  ),
13  # write the abstract and title
14  Abstract(),
15  Title(),
16)
17# run the graph
18paper_result = expression(’Write a scientific paper’)
19# access linker to retreive the results from the method expression
20method = expr.linker.find(’Method’)
21# print result of the method expression
22print(method)
Figure 10: We illustrate the computational graph for the Paper expression. Each node represents an instance of an expression with distinct properties and behaviors. The edges denote the reference relationship between expressions. The blue highlighted nodes mark the main sequence of expressions, used to create the paper, such as Method, RelatedWork, Abstract, etc.
@@ Appendix D Installation
This command will install the latest version of SymbolicAI along with its core dependencies, enabling the integration of the framework into Python applications. If you intend to use the framework with local engines444 The local engines are experimental and are run on your local machine. For more details, refer to the ”Local Neuro-Symbolic Engine” section in the documentation., or with engines powered by external APIs such as OpenAI’s API, additional installation steps are required.
@@ D.1 Engine Configuration
After setting up the API keys, the SymbolicAI library is imported in Python using:
@@ D.2 Optional Installations
D.3 Benchmark
To run the benchmark evaluation you first need to install the ExtensityAI/benchmark plugin:
1sympkg i ExtensityAI/benchmark
Then you can go to the \$HOME/.symai/packages/ExtensityAI/benchmark directory and run the benchmark through:
1python test.py
To compile models optimized for your machine, follow the instructions from the ExtensityAI/benchmark repository.
@@ E.1 Fuzzy Comparison
For instance, let’s consider the use of fuzzy555 Not related to fuzzy logic, which is a topic under active consideration. comparisons. Within SymbolicAI, it enables more adaptable and context-aware evaluations, accommodating the inherent uncertainties and variances often encountered in real-world data.
@@ E.2 Dynamic Casting
E.3 Translation

### v3

@@ 7 Evaluation
Figure 6: We evaluate GPT-4 Turbo, GPT-3.5 Turbo, Gemini-1.0 Pro, LLaMA2-Chat 13B, LLaMA3-Chat 8B, LLaMA3-Chat 70B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs). We denote the VERTEX scores for each category as a normalized value between 00 and $1$, where higher values are better. The VERTEX score is measured according to a reference baseline and normalized by random sequences to exclude noise and similarities among references distributions to rescale solutions. The shown scores are an average over all tests per category and across 8 different seeds per test.
Benchmarks GPT-4 Turbo GPT-3.5 Turbo Gemini 1.0 Pro LLaMA 2 13B LLaMA 3 8B LLaMA 3 70B Mistral 7B Zephyr 7B Random Associations 0.94 0.51 0.93 0.40 0.46 0.57 0.46 0.27 0.08 Modality 0.83 0.65 0.81 0.50 0.43 0.63 0.39 0.43 0.07 Code 0.75 0.63 0.61 0.13 0.25 0.30 0.19 0.13 0.00 Logic 0.48 0.46 0.28 0.46 0.21 0.11 0.53 0.47 0.00 Graphs 0.39 0.34 0.26 0.06 0.05 0.15 0.03 0.03 0.00 Total 0.68 0.52 0.58 0.31 0.28 0.35 0.32 0.27 0.03
@@ 8 Discussion
In this section, we address the evaluation results, auxiliary findings and limitations of SymbolicAI and the future directions we are focusing on. Some of the limitations stem from the inherent constraints of current technologies and dependencies on third-party systems. Additionally, the nuanced complexities of working with generative models presents further challenges.
8.1 Results
In Figure 6 we show the VERTEX score for all five evaluation categories on 8 different state-of-the-art models. We show the aggregated results per category, meaning the average score among all tests averaged per category and average across 8 different seeds per test. The VERTEX score is normalized between 00 and $1$, where higher values are better. Our score is non-linear due to its nature of using non-linear kernels, and captures semantic, ordinal and relative structures among the data samples. However, since our score is highly dependent on the quality of the underlying embedding model, it may omit to capture fine-grained syntactic differences such as ‘Hello‘ vs ‘hello‘.
In our experiments, we have noticed that for associative predictions and multi-modal bindings, GPT-4 Turbo is on par with Gemini-1.0 Pro. Furthermore, there is still a large gap between open-source contestants such as LLaMA 3 even with 70B parameters compared to the closed-source alternatives from OpenAI and Google. For the rest of the experiments, we see that GPT-4 almost always outperforms all other models, except for the functional logic components category. Here, we analyzed results and found that the larger models sometimes take shortcuts by automatically returning the solution and answering that the task instructions are too complex for such a straight-forward puzzle query. However, we would rather state in general that for logic-based, planning and scheduling tasks all models act unreliably, even if slight performance differences between the models are seen in the plot. We believe this is in part due to lack of training data specifically for workflows, planning and scheduling tasks, and to imprecision in generating reliably structured output formats, such as custom DSLs or other custom in-context instructed formats. This also stems from their instruction fine-tuning, since most models are chat-based models and offer verbose responses which need to be suppressed or post-processed.
We see similar performance between GPT-3.5 Turbo and LLaMA 3 70B except for the logical and graphs evaluations. We found that LLaMA 3 70B has a tendency to ask questions back if it does not understand the request instead of following the specified instructions provided. We assume this also stems from the chat-based instruction fine-tuning. Zephyr 7B and Mistral 7B have shown on par capabilities in functional logic components with larger models, however fail in program synthesis and hierarchical computational graphs experiments. We observe that they perform well when resolving the overloaded logic operators such as OR, AND and XOR, and show decent performance for text generation, but fail to resolve more complex instructions.
8.2 Limitations
Framework
Since the framework interfaces with many tools and API services, it requires a substantial engineering feat to integrate all available functionalities and keep the API-based services up-to-date. For us this means, that although we support a variety of tools and frameworks like Selenium, WolframAlpha, or Z3, we only scratch the surface of these tools. Moreover, the utilization of grammar-based constraints validations is still experimental and limited in functionality for specific formats such as JSON and HTML. Finally, we encounter also challenges related to engineering parallelization and multiprocessing of prompts, since the concurrent execution is non-trivial, especially with intricacies of Python process management.
@@ 8.2 Limitations
Our empirical measure is limited by the expressiveness of the embedding model and how well it captures the nuances in similarities between two representations. Furthermore, the obtained similarity scores are highly non-linear and difficult to interpret. For instance, two representations may address the same topic, such as the problem description and its respective solution, however, when measuring their similarity we obtain similarity scores of $\sim 70\%$. We normalize this by subtracting an inherent baseline and randomness effect, however, to ensure a more holistic and robust measurement we would need a significantly larger amount of baselines and experiments. Since we were very limited in the availability of development resources, and some presented models are only addressable through costly API walls. We are actively seeking sponsors to scale our solution and offer a more compelling benchmark suite in the future.
@@
This research is also limited by current LLM’s capacity for reasoning and generalization. Although progress has been made, models are still prone to hallucinations and reasoning errors, especially when dealing with abstract, novel, or highly complex problem statements (Marcus, 2020). Furthermore, our framework’s reliance on the model’s ability to grasp the semantics of operations can be influenced by the training data and the model’s innate biases and conceptual understanding (Mahowald et al., 2023). We also point out that the initial development of SymbolicAI started with the GPT family of models, and we may encounter innate biases in prompt design and expressiveness when utilizing other reference models. However, we also point out that prompt engineering instruction-based statements is not a reliable direction for improvement. We instead advocate for enhancing the resilience of models through fault tolerance, focusing on their ability to better follow semantic instructions, not syntactic idiosyncrasies. Another concern is how to assess the disentanglement of evaluations of models on downstream tasks, to avoid testing on training samples, especially for closed-source solutions like GPT.
@@ 8.3 Future Work
The goal for Algorithm 1 is to be utilized by an advanced learning agent. This agent, employing reinforcement learning methodologies (Ouyang et al., 2022; Li et al., 2023; Rafailov et al., 2023), could leverage our evaluation measure in existing benchmarks (Milani et al., 2020; Swazinna et al., 2022; Schweighofer et al., 2022) as a means to obtain reward signals to addresses a central problem in reinforcement learning, namely credit assignment (Sutton, 1984; Arjona-Medina et al., 2019; Holzleitner et al., 2020; Patil et al., 2020; Widrich et al., 2021; Dinu et al., 2022). Over time, it aims to develop the ability to autonomously generate its own plans, efficiently schedule tasks and subtasks, and carefully select the most suitable tools for each task. Our protocol lays the groundwork for this agent to learn and expand its base set of capabilities (Amaro et al., 2023), moving towards more sophisticated, self-referential orchestration of multi-step tasks. We’ve already noticed that research is shifting towards this type of methodology (Yuan et al., 2024). Furthermore, in Section 7 we’ve only considered a sequential scheduler. However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution. In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a; b; Lahlou et al., 2023). Lastly, we also discuss limitations in Appendix Section 8.2 with further opportunities for future improvements.
9 Conclusion
@@
In this work, we introduced SymbolicAI, a framework that unifies generative models with an array of solvers, blending the strengths of symbolic and sub-symbolic AI paradigms within a cohesive NeSy framework. SymbolicAI equips researchers and practitioners with a comprehensive toolkit to develop contextualized and explainable NeSy AI systems capable of addressing diverse challenges effectively. We also introduce a quality measure and a benchmark for comparing and evaluating a wide range of computational tasks. SymbolicAI provides a basis for further research in advanced program synthesis, hierarchical computational graphs, the development of self-referential systems, and the integration of probabilistic methods with AI design for creating autonomous agents.
@@ Acknowledgement
Appendix A Broad AI and Neuro-Symbolic Systems
Our work focuses on broad AI (Hochreiter, 2022) (see Figure 7) through the integration of symbolic and sub-symbolic AI methodologies. Broad AI extends beyond restricted focus on single-task performance of narrow AI. In broad AI, systems are engineered to handle a range of tasks with a high degree of autonomy, utilizing sensory input, accumulated experiences, and previously developed skills.
NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023). This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.
A.1 Broader Impact
With LLMs becoming more and more accessible, progress recently made possible by the vast open source contributions from Köpf et al. (2023); Touvron et al. (2023); Taori et al. (2023); Xu et al. (2023); Geng et al. (2023); Biderman et al. (2023), embedded accelerators for LLMs — or more generally NeSY engines — will be ubiquitous in future computation platforms, such as wearables, smartphones, tablets, consoles, or notebooks. Although current execution cycles are slow and error-prone, we expect to see further performance gains through improved operating system level optimizations, dedicated GPU-centric hardware refinement, and improved software interoperability. We believe that modern programming paradigms should natively support probabilistic concepts and provide a boilerplate-free set of features for constructing and evaluating generative computational graphs. This includes but is not limited to compositional, parallelizable, and simulation-based executions with polymorphic operations and self-referential structures. Current programming languages often have disjointed or makeshift solutions for these concepts in the context of generative processes. We believe integral probabilistic support for these concepts into modern software and hardware will unlock new programming paradigms that can fully take advantage of generative architectures. We hope the community will consider these ideas as essential components of contemporary computing.
Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure. Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs. This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time. We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023). We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section A) and is a requirement for developing artificial general intelligent agents.
Appendix B Connection between Fréchet Distance and Maximum Mean Discrepancy
@@
In the next section, we want to further elaborate on the MMD and a possible score, that can be derived from it.
B.1 Extended Simplification of the MMD Calculation
@@
Appendix C Structure
@@
Type inheritance in SymbolicAI is leveraged to create new expressions, which are specialized forms of Symbol objects designed to represent parts of a computational graph. Expressions extend the capabilities of Symbol objects by providing a structured way to define complex functionalities that can later be evaluated to produce new Symbol objects or modify existing ones. In SymbolicAI, an Expression object inherits the properties of Symbol objects while also being able to define its own unique behavior through a forward method, which is analogous to a computational graph node’s evaluation function. Figure 10 gives an overview of an exemplary inheritance branch. Each Expression must feature a forward method, which must be overwritten to define its behavior. The inherited __call__ method invokes the forward method, evaluating the expression and returning the result. This design pattern facilitates lazy evaluation of expressions, allowing for complex composition of expressions.
@@
Decorators brings several advantages (Beazley, 2009; Martelli et al., 2005; Summerfield, 2010; Lutz, 2013):
@@
  • Adaptability: Through decorators we can easily modify or extend the behavior of operations without changing their core implementation. This flexibility facilitates the generation of adaptive workflows and reliable fallback mechanisms when experimental implementations do not fulfill required constraints.
@@
In practice, consider the Expression class, which extends the functionality of the Symbol class. When composing a Sequence of Expression objects, we are effectively composing operations in a predetermined order.
For instance, an expression like:
@@ Appendix D Installation
This command will install the latest version of SymbolicAI along with its core dependencies, enabling the integration of the framework into Python applications. If you intend to utilize the framework with local engines555 The local engines are experimental and are run on your local machine. For more details, refer to the ”Local Neuro-Symbolic Engine” section in the documentation., or with engines powered by external APIs such as OpenAI’s API, additional installation steps are required.
@@ D.1 Engine Configuration
After setting up the API keys, the SymbolicAI library is imported in Python as following:
@@ E.1 Fuzzy Comparison
For instance, let’s consider fuzzy666 Not related to fuzzy logic, which is a topic under active consideration. comparisons. Within SymbolicAI, it enables more adaptable and context-aware evaluations, accommodating the inherent uncertainties and variances often encountered in real-world data.
@@ E.2 Dynamic Casting
E.3 Symbols and Embeddings
It is worth noting that encoding a complex object into a string sometimes precludes the object reconstitution. However, this concern does not substantially impede our methodology as we can employ approximations or proxy representations stored by the vector-valued property to effectively re-map objects. These representations are obtained through respective embedding models.
E.4 Translation